Given the population 5, 10, 15, 20, 25 a) How many samples of size 3, can be drawn with replacement from this population b) Compute and tabulate the sampling distribution of the mean from samples of size 3. c) Verify the results of mean and variance of sampling distribution of the mean. 


1
Expert's answer
2021-05-28T10:20:41-0400

Mean



"\mu=\dfrac{5+10+15+20+25}{5}=15"

Variance


"\sigma^2=\dfrac{1}{5}\big((5-15)^2+(10-15)^2+(15-15)^2""(20-15)^2+(25-15)^2\big)=50"


Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{50}=5\sqrt{2}\approx7.071068"


We have population values "5,10,15,20,25" population size "N=5" and sample size "n=3." Thus, the number of possible samples which can be drawn with replacement is


"n^r=5^3=125""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 5,5, 5 & 5 \\
 \hdashline
 2 & 5,5,10 & 20/3 \\
 \hdashline
 3 & 5,5,15 & 25/3 \\
 \hdashline
 4 & 5,5,20 & 10 \\
 \hdashline
 5 & 5,5,25 & 35/3 \\
 \hline
6 & 5,10,5 & 20/3 \\
 \hline
7 & 5,10,10 & 25/3 \\
 \hline
8 & 5,10,15 & 10 \\
 \hline
9 & 5,10,20 & 35/3 \\
 \hline
10 & 5,10,25 & 40/3 \\
 \hline
11 & 5,15,5 & 25/3 \\
 \hline
12 & 5,15,10 & 10 \\
 \hline
13 & 5,15,15 & 35/3 \\
 \hline
14 & 5,15,20 & 40/3 \\
 \hline
15 & 5,15,25 & 15 \\
 \hline
16 & 5,20,5 & 10 \\
 \hline
17 & 5,20,10 & 35/3 \\
 \hline
18 & 5,20,15 & 40/3 \\
 \hline
19 & 5,20,20 & 15 \\
 \hline
20 & 5,20,25 & 50/3 \\
 \hline
21 & 5,25,5 & 35/3\\
 \hline
22 & 5,25,10 & 40/3 \\
 \hline
23 & 5,25,15 & 15 \\
 \hline
24 & 5,25,20 & 50/3 \\
 \hline
25 & 5,25,25 & 55/3 \\
 \hline
26 & 10,5, 5 & 20/3 \\
 \hdashline
 27 & 10,5,10 & 25/3 \\
 \hdashline
 28 & 10,5,15 & 10 \\
 \hdashline
 29 & 10,5,20 & 35/3 \\
 \hdashline
 30 & 10,5,25 & 40/3 \\
 \hline
31 & 10,10,5 & 25/3 \\
 \hline
32 & 10,10,10 & 10 \\
 \hline
33 & 10,10,15 & 35/3 \\
 \hline
34 & 10,10,20 & 40/3 \\
 \hline
35 & 10,10,25 & 15 \\
 \hline
36 & 10,15,5 & 10 \\
 \hline
37 & 10,15,10 & 35/3\\
 \hline
38 & 10,15,15 & 40/3 \\
 \hline
39 & 10,15,20 & 15 \\
 \hline
40 & 10,15,25 & 50/3 \\
 \hline
41 & 10,20,5 & 35/3 \\
 \hline
42 & 10,20,10 & 40/3 \\
 \hline
43 & 10,20,15 & 15 \\
 \hline
44 & 10,20,20 & 50/3 \\
 \hline
45 & 10,20,25 & 55/3 \\
 \hline
46 & 10,25,5 & 40/3\\
 \hline
47 & 10,25,10 & 15 \\
 \hline
48 & 10,25,15 & 50/3 \\
 \hline
49 & 10,25,20 & 55/3 \\
 \hline
50 & 10,25,25 & 20 \\
 \hline
 51 & 15,5, 5 & 25/3 \\
 \hdashline
 52 & 15,5,10 & 10 \\
 \hdashline
 53 & 15,5,15 & 35/3 \\
 \hdashline
 54 & 15,5,20 & 40/3 \\
 \hdashline
 55 & 15,5,25 & 15 \\
 \hline
56 & 15,10,5 & 10 \\
 \hline
57 & 15,10,10 &35/3 \\
 \hline
58 & 15,10,15 & 40/3 \\
 \hline
59 & 15,10,20 & 15 \\
 \hline
60 & 15,10,25 & 50/3 \\
 \hline
61 & 15,15,5 & 35/3 \\
 \hline
62 & 15,15,10 & 40/3 \\
 \hline
63 & 15,15,15 & 15 \\
 \hline
64 & 15,15,20 & 50/3 \\
 \hline
65 & 15,15,25 & 55/3 \\
 \hline
66 & 15,20,5 & 40/3 \\
 \hline
67 & 15,20,10 & 15 \\
 \hline
68 & 15,20,15 & 50/3 \\
 \hline
69 & 15,20,20 & 55/3 \\
 \hline
70 & 15,20,25 & 20 \\
 \hline
71 & 15,25,5 & 15\\
 \hline
72 & 15,25,10 & 50/3 \\
 \hline
73 & 15,25,15 & 55/3 \\
 \hline
74 & 15,25,20 & 20 \\
 \hline
75 & 15,25,25 & 65/3 \\
 \hline
76 & 20,5, 5 & 10 \\
 \hdashline
 77 & 20,5,10 & 35/3 \\
 \hdashline
 78 & 20,5,15 & 40/3 \\
 \hdashline
 79 & 20,5,20 & 15 \\
 \hdashline
 80 & 20,5,25 & 50/3 \\
 \hline
81 & 20,10,5 & 35/3 \\
 \hline
82 & 20,10,10 & 40/3 \\
 \hline
83 & 20,10,15 & 15 \\
 \hline
84 & 20,10,20 & 50/3 \\
 \hline
85 & 20,10,25 & 55/3 \\
 \hline
86 & 20,15,5 & 40/3 \\
 \hline
87 & 20,15,10 & 15\\
 \hline
88 & 20,15,15 & 50/3 \\
 \hline
89 & 20,15,20 & 55/3 \\
 \hline
90 & 20,15,25 & 20 \\
 \hline
91 & 20,20,5 & 15 \\
 \hline
92 & 20,20,10 & 50/3 \\
 \hline
93 & 20,20,15 & 55/3 \\
 \hline
94 & 20,20,20 & 20 \\
 \hline
95 & 20,20,25 & 65/3 \\
 \hline
96 & 20,25,5 & 50/3\\
 \hline
97 & 20,25,10 & 55/3 \\
 \hline
98 & 20,25,15 & 20 \\
 \hline
99 & 20,25,20 & 65/3 \\
 \hline
100 & 20,25,25 & 70/3 \\
 \hline
 101 & 25,5, 5 & 35/3 \\
 \hdashline
 102 & 25,5,10 & 40/3 \\
 \hdashline
 103 & 25,5,15 & 15 \\
 \hdashline
 104 & 25,5,20 & 50/3 \\
 \hdashline
 105 & 25,5,25 & 55/3 \\
 \hline
106 & 25,10,5 & 40/3 \\
 \hline
107 & 25,10,10 & 15 \\
 \hline
108 & 25,10,15 & 50/3 \\
 \hline
109 & 25,10,20 & 55/3 \\
 \hline
110 & 25,10,25 & 20 \\
 \hline
111 & 25,15,5 & 15 \\
 \hline
112 & 25,15,10 & 50/3 \\
 \hline
113 & 25,15,15 & 55/3 \\
 \hline
114 & 25,15,20 & 20 \\
 \hline
115 & 25,15,25 & 65/3 \\
 \hline
116 & 25,20,5 & 50/3 \\
 \hline
117 & 25,20,10 & 55/3 \\
 \hline
118 & 25,20,15 & 20 \\
 \hline
119 & 25,20,20 & 65/3 \\
 \hline
120 & 25,20,25 & 70/3 \\
 \hline
121 & 25,25,5 & 55/3\\
 \hline
122 & 25,25,10 & 20 \\
 \hline
123 & 25,25,15 & 65/3 \\
 \hline
124 & 25,25,20 & 70/3 \\
 \hline
125 & 25,25,25 & 25 \\
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 15/3 & 1& 1/125 & 3/75 & 9/45 \\
 \hdashline
 20/3 & 3 & 3/125 & 12/75 & 48/45 \\
 \hdashline
 25/3 & 6 & 6/125 & 30/75 & 150/45 \\
 \hdashline
 30/3 & 10 & 10/125 & 60/75 & 360/45 \\
 \hdashline
 35/3 & 15 & 15/125 & 105/75& 735/45 \\
 \hdashline
40/3 & 18 & 18/125 & 144/75 & 1152/45 \\
 \hdashline
 45/3 & 19 & 19/125 & 171/75 & 1539/45 \\
 \hdashline
 50/3 & 18 & 18/125 & 180/75 & 1800/45 \\
 \hdashline
 55/3 & 15 & 15/125 & 165/75 & 1815/45 \\
 \hdashline
 60/3 & 10 & 10/125 & 120/75 & 1440/45 \\
 \hdashline
 65/3 & 6 & 6/125 & 78/75 & 1014/45 \\
 \hdashline
 70/3 & 3 & 3/125 & 42/75 & 588/45 \\
 \hdashline
 75/3 & 1& 1/125 & 15/75 & 225/45 \\
 \hdashline
 
 Total & 125 & 1 & 225/15 & 725/3 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{225}{15}=15"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=15=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{725}{3}-(15)^2=\dfrac{50}{3}"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{50}{3}}\approx4.082483"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}=\dfrac{50}{3},True"